OpenStudy (anonymous):

What are the zeroes of f(x) = x^2 − 6x + 8? x = −4, 2 x = −4, −2 x = 4, 2 x = 4, −2

OpenStudy (anonymous):

I think it is C.

OpenStudy (sohailiftikhar):

yes it's c

OpenStudy (anonymous):

Thank you!

OpenStudy (sohailiftikhar):

np :)

OpenStudy (anonymous):

Can you help me with another question?

OpenStudy (sohailiftikhar):

ok

OpenStudy (anonymous):

Give me 1 min :]

OpenStudy (sohailiftikhar):

ok

OpenStudy (anonymous):

@sohailiftikhar Part A The first part of Ray and Kelsey's roller coaster is a curved pattern that can be represented by a polynomial function. Ray and Kelsey are working to graph a third-degree polynomial function that represents the first pattern in the coaster plan. Ray says the third-degree polynomial has 4 intercepts. Kelsey argues the function can have as many as 3 zeros only. Is there a way for the both of them to be correct? Explain your answer. Kelsey has a list of possible functions. Pick one of the g(x) functions below and then describe to Kelsey the key features of g(x), including the end behavior, y-intercept, and zeros. g(x) = x3 − x2 − 4x + 4 g(x) = x3 + 2x2 − 9x − 18 g(x) = x3 − 3x2 − 4x + 12 g(x) = x3 + 2x2 − 25x − 50 g(x) = 2x3 + 14x2 − 2x − 14 Create a graph of the polynomial function you selected from Question 2. Part B The second part of the new coaster is a parabola. Ray needs help creating the second part of the coaster. Create a unique parabola in the pattern f(x) = ax2 + bx + c. Describe the direction of the parabola and determine the y-intercept and zeros. The safety inspector notes that Ray also needs to plan for a vertical ladder through the center of the coaster's parabolic shape for access to the coaster to perform safety repairs. Find the vertex and the equation for the axis of symmetry of the parabola, showing your work, so Ray can include it in his coaster plan. Create a graph of the polynomial function you created in Question 4. Part C Now that the curve pieces are determined, use those pieces as sections of a complete coaster. By hand or by using a drawing program, sketch a design of Ray and Kelsey's coaster that includes the shape of the g(x) and f(x) functions that you chose in the Parts A and B. You do not have to include the coordinate plane. You may arrange the functions in any order you choose, but label each section of the graph with the corresponding function for your instructor to view.

OpenStudy (anonymous):

@sohailiftikhar

OpenStudy (anonymous):

@sohailiftikhar